### v2

# Davide Frizzo

Abstract
Anomaly Detection (AD) is crucial in industrial settings to streamline operations by detecting underlying issues. Conventional methods merely label observations as normal or anomalous, lacking crucial insights. In Industry 5.0, interpretable outcomes become desirable to enable users to understand the rational under model decisions. This paper presents the first industrial application of ExIFFI, a recent approach for fast, efficient explanations for the Extended Isolation Forest (EIF) AD method. ExIFFI is tested on three industrial datasets, demonstrating superior explanation effectiveness and computational efficiency compared to other state-of-the-art explainable AD models.
Keywords Anomaly Detection, Explainable Artificial Intelligence, Industrial Internet of Things
AI
Artificial Intelligence
AD
Anomaly Detection
IF
Isolation Forest
EIF
Extended Isolation Forest
DIFFI
Depth Based Isolation Forest Feature Importance
AcME-AD
Accelerated Model Explanations for Anomaly Detection
ML
Machine Learning
IoT
Internet of Things
IIoT
Industrial Internet of Things
XAI
Explainable Artificial Intelligence
SHAP
SHapley Additive exPlanations
GFI
Global Feature Importance
LFI
Local Feature Importance
PIADE
Packaging Industry Anomaly DEtection
TE
Tennessee Eastman
TEP
Tennessee Eastman Process
LODA
Lightweight on-line detector of anomalies
1 Introduction
In recent decades, the rapid expansion of Machine Learning (ML) applications has significantly propelled advancements in Internet of Things (IoT) technologies, leading to the widespread deployment of interconnected sensor networks [1]. These developments are particularly notable in the context of Industrial Internet of Things (IIoT), which focuses on real-time monitoring of industrial assets to optimize maintenance and enhance production efficiency[2]. As these IoT systems generate vast datasets, ML algorithms play a crucial role in analyzing this data to extract predictive insights, underscoring a symbiotic relationship between ML and IIoT that mutually enhance their capabilities [3].
A key application within this context is unsupervised AD [4], especially prominent in settings where labeling data is impractical. Among AD methods, models based on Isolation Forest (IF) [5], in particular EIF [6] stands out due to its speed, low memory requirements, and high performance, making it valuable in industrial settings for swiftly pinpointing anomalies with exceptional performances [7]. However, the challenge extends beyond merely detecting anomalies; understanding their causes is essential for effective resolution [8]. Here, Explainable Artificial Intelligence (XAI) becomes crucial by making the outputs of ML models like EIF transparent and actionable, thus supporting informed decision-making and enhancing system reliability within IIoT environments [9].
This paper specifically evaluates the Extended Isolation Forest Feature Importance (ExIFFI) algorithm, developed by Arcudi et al.[10], which provides a time-efficient and tailored interpretative approach to the EIF model. Moreover, [10] additionally introduces a modification of the EIF algorithm, named Extended Isolation Forest Plus ( $\text{EIF}^{+}$) which improves the generalization performances of EIF and can also be interpreted by ExIFFI. The algorithm’s effectiveness is showcased through its application on real world IIoT datasets, demonstrating how it enhances the decision-making process in industrial settings.
This paper is the extension of [11]. The improvements over the original version consist of:
  • An evaluation of the explained models in terms of AD performances, detailed in 4.3
  • Additional images to showcase the results (i.e. Figures 3 , 4 , 6)
  • The addition of a new benchmark dataset, CoffeeData , introduced in 4.1.3
The paper is organized as follows: in Section 2 ExIFFI is contextualized within Industry 5.0, Section 3 explains the operational mechanics of the algorithm, Section 4 presents the benchmark datasets and the application of ExIFFI on them. Finally, Section 5 summarizes the key findings and discusses potential directions for future research.
2 Related Work
The progression from Industry 4.0 [12] to Industry 5.0 [13] marks a shift from a focus on enhancing operational performance through automation and data exchange, primarily using cyber-physical systems and IoT, to emphasizing human-centric outcomes [14]. Industry 5.0 integrates human creativity with the technical prowess of Artificial Intelligence (AI) and robotics, highlighting the need for transparency and interpretability in machine learning models to build trust and aid decision-making.
This need is addressed by implementing XAI in various industrial applications, including fault detection in machinery [15], process monitoring across sectors from semiconductors to home appliances [16, 17, 18], and predictive maintenance [19]. The ExIFFI interpretation algorithm of the EIF model used for AD exemplifies this trend. It offers a fast, precise interpretation method crucial for IIoT environments, where performance and user-centric understanding are imperative for efficiency and accessibility in modern manufacturing setups.
In this paper ExIFFI is evaluated against various interpretation techniques, including both ad-hoc  [^1] Ad-hoc interpretation algorithms are built into models for inherent transparency. Post-hoc methods clarify complex model decisions after development but can be computationally intensive and approximate. [20] methods like Depth Based Isolation Forest Feature Importance (DIFFI) [21], tailored for the Isolation Forest, and post-hoc methods such as KernelSHAP [22] and Accelerated Model Explanations for Anomaly Detection (AcME-AD) [23] [24].
KernelSHAP is notable for its ability to interpret diverse models through a model-agnostic, kernel-based approximation of SHapley Additive exPlanations (SHAP) values. However, it can be computationally demanding, particularly with large datasets. In contrast, AcME-AD streamlines the explanation process, making it ideal for scenarios that require swift interpretability compared to KernelSHAP.
Although ExIFFI and DIFFI are less flexible compared to post-hoc methods, they significantly boost computational speed, delivering reliable interpretations up to 100 times faster than post-hoc models 4. Additionally, ExIFFI leverages the advanced structure of the EIF, that outperforms the traditional IF [7] by avoiding its structural biases [6]. This enables ExIFFI to assess feature significance both efficiently and accurately.
3 Proposed Approach
The ExIFFI, similarly to the DIFFI with the IF, leverages the structure of the forest generated by the EIF to assess the significance of each feature in determining whether a sample is anomalous.
In the structure of the EIF, each tree $t$ in the forest $\mathcal{T}$ consists of nodes that partition the space using a hyperplane $\mathcal{H}_{k}^{t}$. Each hyperplane is defined by a normal vector $\mathbf{v}_{k}^{t}$ and an intercept point $\mathbf{p}^{t}_{k}$. The hyperplane splits the data sample $X_{k}^{t}$ into two distinct subsets $L^{t}_{k}$ and $R^{t}_{k}$, with $L^{t}_{k}\cup R^{t}_{k}=X_{k}^{t}$ and $L^{t}_{k}\cap R^{t}_{k}=\varnothing$. The ExIFFI algorithm assesses the feature importance by calculating the imbalance generated by each node for a given sample $x$ as follows:
$\mathbf{\lambda}^{t}_{k}(\mathbf{x})=\begin{cases}\left(\frac{|X^{t}_{k}|}{|L^{t}_{k}|}\right)\text{abs}(\mathbf{v}^{t}_{k}),&\text{if }\mathbf{v}^{t}_{k}\cdot\mathbf{x}>\mathbf{v}^{t}_{k}\cdot\mathbf{p}^{t}_{k}\\ \left(\frac{|X^{t}_{k}|}{|R^{t}_{k}|}\right)\text{abs}(\mathbf{v}^{t}_{k}),&\text{otherwise}\end{cases}$
The feature importance for a sample $x$ within a specific tree $t$ is determined by the summation of importance vectors from all nodes $k$ that $x$ traverses on its path to the leaf: $\mathbf{I}_{t}(x)=\sum_{k\in\mathcal{P}^{t}_{x}}\mathbf{\lambda}_{k}^{t}(x)$. The overall feature importance for $x$ across the entire forest is then computed by aggregating these importance vectors from all the trees that $x$ encounters: $\mathbf{I}(x)=\sum_{t\in\mathcal{T}}\mathbf{I}_{t}(x)$.
This importance vector $\mathbf{I}(x)$ assesses the influence of each feature on the isolation of $x$, identifying the ones predominantly contributing to the classification $x$ as an anomaly. To correct for potential biases caused by features being sampled more frequently, $\mathbf{I}(x)$ is normalized by the sum of vectors orthogonal to the hyperplanes of the nodes that $x$ traverses in each tree, denoted as $\mathbf{V}(x)$: $\mathbf{V}(x)=\sum_{t\in\mathcal{T}}\sum_{k\in\mathcal{P}^{t}_{x}}\mathbf{v}_{k}^{t}$.
Grouping together the importance vectors assigned to all the samples in the input dataset, ExIFFI is equipped with two possible outputs: Global Feature Importance (GFI) and Local Feature Importance (LFI). GFI returns a single vector which associates a score to each feature, quantifying its overall importance in discriminating between inliers and outliers. It can be obtained as follows: $\mathbf{GFI}=\frac{\hat{\mathbf{I}}_{O}}{\hat{\mathbf{I}}_{I}}$, where $\hat{\mathbf{I}}_{O}$ and $\hat{\mathbf{I}}_{I}$ are the importance vectors computed over the set of outliers and inliers respectively. On the other hand, the Local Feature Importance ($\mathbf{LFI}$) for $x$ is calculated as: $\mathbf{LFI}(x)=\frac{\mathbf{I}(x)}{\mathbf{V}(x)}$. This metric offers a refined and normalized measure of which features significantly impact the classification of a single sample $x$ as anomalous, thereby enhancing the interpretability of anomaly detection models and facilitating targeted interventions based on the critical features identified.
4 Experimental Results
This section presents the results of applying ExIFFI to two publicly accessible and a private dataset derived from industrial processes, which serve as benchmarks for evaluating ExIFFI’s effectiveness within real-world contexts. The datasets include Tennessee Eastman Process (TEP), which offers synthetic data with established ground truth for anomaly-inducing features [25], Packaging Industry Anomaly DEtection (PIADE), which encapsulates typical real-world challenges such as unlabeled and high-dimensional data [26] and CoffeeData, a confidential dataset containing sensor measurements recorded during the brewing phase of capsules on coffee machines.
The section is organized as follows: firstly, datasets are presented in 4.1, the experimental setup is provided in 4.2, AD model performances are reported in 4.3, global and local interpretability are assessed in 4.4.1, 4.4.2, 4.6.1 respectively, and four different interpretability models are compared in 4.4.3, 4.6.2 by means of the Feature Selection Proxy Task.
The outcomes here discussed are achieved using Python as the base language to implement the method and C to optimize functions embedded within the performance-critical segments of the Python code. [^2] The source code of this project is available in a public repository, with reproducible results: https://github.com/francesco-borsatti-unipd/ExIFFI_Industrial_Test.
4.1 Industrial IoT datasets
In the following, we provide an in-depth description of the structure of the benchmark datasets used in this study (i.e. TEP, PIADE and CoffeeData) pinpointing their key characteristics and differences.
The datasets exhibit fundamental disparities representing different industrial scenarios. Specifically PIADE comprises machine alarms and statuses, while TEP and CoffeeData are characterized by time-series data. These differences underscore the wide range of signal types on which the presented approach can be applied. Another key difference between the three relies on how data are generated: PIADE and CoffeeData are real datasets collected from existing machines, while TEP was obtained through simulations (albeit precise).
4.1.1 TEP dataset
The TEP dataset [25] collects data coming from the Tennessee Eastman (TE) process, crucial for the production of fibers, chemicals and advanced materials for everyday purpose, as detailed in [27]. TEP data are obtained as the result of simulations of the TE process which can generate normal or faulty runs. Each simulation run produces a time series composed of 500 samples and 52 process variables. Each sample can be categorized into 21 classes: Class 0 denotes normal samples while Class 1-20 represent faulty instances.
Importantly, prior knowledge on the features associated with the first 15 faults has been documented [28], making it possible to assess the correctness of the explanations proposed by ExIFFI and facilitating comparative analysis with alternative interpretative algorithms through the Feature Selection proxy task, as elaborated in Section 4.2.
Moreover, the presence of labeled data provides the possibility of assessing the performance of $\text{EIF}^{+}$, the model interpreted by ExIFFI, as an Anomaly Detection model. $\text{EIF}^{+}$ displays excellent performance in identifying abnormal chemical processes, as detailed in Table 2.
In order to provide a better visualization of the results, we consider a downsampled version of the TEP dataset. Seventy three simulations were selected: 70 normal simulations and 3 faulty ones, for a total of 35600 samples and 52 features. In particular, we focus on fault type IDV12 on which domain knowledge about the most important features is available [28]. The fault considered relates to the condenser cooling water temperature and the root cause feature is xmeas_11, the separator temperature measure.
4.1.2 PIADE dataset
The Alarm logs Packaging equipment dataset (PIADE) dataset [26] is publicly available at  [^3] https://zenodo.org/records/7071747.
This dataset comprises real data produced by packaging machines. Differently from the TEP dataset, PIADE lacks labeled data and there is no documented ground truth on the relevance of the different features used to characterize the industrial processes described. Nevertheless, it is possible to validate the correctness of the provided explanations with domain experts.
Another major difference with respect to the TEP dataset lies in the nature of the encapsulated data. Unlike traditional IIoT datasets composed of sensor measurements, PIADE comprises statistics concerning alarm counts triggered by specific machinery, indirectly encapsulating the normal or anomalous operational states of the equipment.
This data collection strategy is motivated primarily by the substantially lower implementation costs associated with it. Indeed, the alarm logs are directly generated by the packaging machines under investigation without the need of installing additional expensive ad-hoc sensors. This approach has several applications, including Alarm Forecasting [29].
PIADE contains data derived from five machines of identical type but operating under distinct conditions. Consequently, the dataset is partitioned into five sub-datasets, one for each machine. Notably, the experiments described in 4 focused on the second machine (i.e. Equipment_ID equal to 2).
Finally, for what concerns the data structure, alarm logs data are aggregated into 1-hour-long time windows starting from the raw values, resulting in 2725 samples.
The number of features amounts to 162, considering all alarms registered across the five machines. For our experiments, we decided to remove the features connected to alarms and operational state transitions that were never encountered in the equipment considered.
As a result, the final dataset comprises 2725 samples and 71 features.
4.1.3 CoffeeData Dataset
The CoffeeData dataset was firstly introduced in [30]. It is a confidential dataset and consists of multiple time-series representing the brewing process of capsules in coffee machines. The aim of AD in this scenario is to swiftly identify anomalous usages of coffee capsules in order to improve the brewing quality. Anomalous behaviors include using the same capsule twice or more times (i.e. reused capsule) or using the coffee machine without any capsule inside it (i.e. no capsule).
The data were collected through a Yokogawa MX100 device using a sampling frequency of 2 Hz. Eight variables are registered including: heating element voltage and current, water flow rate, temperatures for the heating element bulk, output water, ambient air, water tank, and water pump.
Following the procedure introduced in [30], only the water flow rate feature was considered, as it was the most informative one.
Each time series is trimmed to the first 15 samples of the pre-infusion phase, as this segment provides enough information to address the problem effectively while also allowing for early detection of anomalies, which is critical for practical use.
In order to take into account the temporal dimension embedded in the data, instead of employing tedious and time-consuming handcrafted features as in [30], raw data are used and time stamps are considered as the features which will be ranked by ExIFFI when computing importance scores. Exploiting this strategy is possible to identify the time instants in which anomalies occur. A possible application of this result may be to trigger an alarm at the anomalous time sample to notify the user of an improper usage of the equipment.
Table 1 outlines the dataset composition: standard capsules are part of the inliers group (comprising both original and compatible brands), while there are two kinds of anomalous samples: reused capsules and no capsules.
Table 1: CoffeeData dataset composition.
Brewing Capsule # Samples Class
1st standard 39 Normal
2nd standard 229
any reused capsule 78 Anomaly
no capsule 26
Normal capsules are divided into the first and second brewings categories which represent the situation in which the coffee machine is used for the first time after a long period of rest and thus is cold and the one in which multiple coffees are produced in a short time interval with the machine at a working regime temperature. Clearly, trials of type second brewing are easier to carry out (i.e. there is no need to wait for the machine to cool down) and thus the amount of acquired samples in this condition is higher.
4.2 Experimental Setup
This section provides a description of the techniques used to evaluate and compare the different interpretability methods considered in this study.
4.2.1 Global Importance Assessment
This first experiment leverages the GFI score returned by ExIFFI, described in 3, to rank the attributes in decreasing order of relevance scores in discerning between normal and anomalous samples.
Feature ranking is computed as the average GFI scores over multiple runs, to account for the inherent stochastic nature of ExIFFI model. This Monte Carlo reminiscent approach ensures a robust representation of the distribution of feature importance.
A Score Plot with the top eight ranked features represented as horizontal bars with length proportional to the GFI score is used to depict the results.
Human operators can take advantage from analyzing the Score Plot. The top ranked features are in fact the most likely causes of anomalous behavior which can be exploited by domain experts to take preventive actions and avoid potential failures.
4.2.2 Local Scoremaps
Within an industrial context emphasis is placed on local interpretability, thus the Local Scoremap plot becomes a powerful tool. It focuses on a pair of features comprehending a scatter plot showing the distribution of inliers and outliers and an heatmap to delineate the LFI scores in a grid of points within the considered two-dimensional feature space. The color of each point depends on the feature whose LFI is higher on that point and its intensity is proportional to the relevance difference between the two attributes.
Normally [10], the two most important features are considered to produce the scoremap. However, in case relevance is shared across multiple features the analysis of multiple pairs of attributes is another viable approach.
Different conclusions may be drawn as a consequence. In case outliers are clearly scattered along a single axis, there is likely a single feature dominating the importance scores. On the other hand, relevance may be distributed among a subset of features. In the latter case, anomalies are scattered along an oblique line, and it may be ideal to analyze multiple pairs of features.
4.2.3 Feature Selection Proxy Task
The Feature Selection Proxy task provides a qualitative and quantitative measure of the effectiveness of an interpretation algorithm. This evaluation is based on the following fact: starting from the GFI score ranking, if an interpretation model ranked features aligning to the ground truth the Average Precision of the interpreted model is relatively stable as irrelevant features are progressively eliminated from the feature space. Conversely, the removal of relevant attributes should not increase the Average Precision. The Feature Selection plot thus displays the trend of Average Precision as features are incrementally discarded, in increasing (i.e. inverse approach) and decreasing (i.e. direct approach) order of importance. A random approach (i.e. features dropped in random order) is also included as a baseline.
The Feature Selection proxy task requires access to the AD model performances (i.e. Average Precision) through labelled data. As a consequence this analysis was performed solely on the TEP and CoffeeData datasets.
4.3 AD Performance Comparison
In this section IF, EIF and $\text{EIF}^{+}$ AD models are compared with typical AD metrics for the labeled datasets (i.e. TEP and CoffeeData). Tables 2, 3 report the test set metrics. This evaluation is a crucial step for the application of XAI methods like ExIFFI since their effectiveness highly relies on the accuracy of the explained model.
Table 2: TEP dataset AD metrics for the considered models. Best performance for each metric in bold.
Model
Average
Precision
Precision
ROC
AUC
Fit
Time [s]
Predict
Time [s]
IF 0.90 0.85 0.92 1.90 1.44e-4
EIF 0.93 0.90 0.95 1.25 8.00e-5
$\text{EIF}^{+}$ 0.93 0.91 0.95 1.71 5.14e-5
Table 2 shows how EIF and $\text{EIF}^{+}$ perform better than IF thanks to their higher-complexity data partitioning scheme, leveraging multidimensional splits as opposed to IF.
CoffeeData comprises two evaluation datasets: D1, which contains both first and second brewings and both kinds of anomalies (i.e. no capsule and reused capsule); while D2 is more limited, but equivalent to the one used in previous work [30] (i.e. only second brewing data and no capsule anomalies).
Table 3: CoffeeData dataset AD metrics for the considered models. Best performance for each metric in bold.
Dataset Model
Average
Precision
Precision
ROC
AUC
Fit
Time [s]
Predict
Time [s]
D1 IF 0.74 0.65 0.63 1.18 5.48e-5
EIF 0.77 0.66 0.63 0.81 3.65e-5
$\text{EIF}^{+}$ 0.79 0.65 0.64 0.99 4.87e-5
D2 IF 0.91 0.96 0.97 0.96 3.04e-5
EIF 0.97 0.89 0.98 0.74 5.48e-5
$\text{EIF}^{+}$ 0.99 0.96 0.99 0.96 4.26e-5
AD model performances are showcased in Table 3. The increased difficulty associated with D1 translates in AD metrics that are slightly worse than the ones observed on TEP and PIADE datasets, but still the best performing model is $\text{EIF}^{+}$ followed by EIF and IF. For what concerns D2 metrics are on par with the ones showcased in [30] where the model was exploiting hand crafted features while raw time samples are used in this setting.
4.4 Case Study I: TEP Dataset
4.4.1 Global Feature Importance
Figure 1 exhibits the Score Plot for the TEP dataset. Features xmeas_22 and xmeas_11 emerge as more relevant than others, aligning with the ground truth [28] since xmeas_11 is the root cause for fault IDV12. Furthermore, the Sign Directed Graph in [28] proves also the saliency of xmeas_12 (i.e. the Separator cooling water outler temperature), a direct consequence of xmeas_11. The causal relationship between these two attributes leads the model to position them on top of the GFI ranking.
Figure 1: GFI Score Plot TEP dataset.
4.4.2 Local Interpretability
Figure 2 groups together the Local Scoremaps produced by ExIFFI for three distinct pairs of attributes within the TEP dataset. This presentation facilitates a comprehensive analysis of how anomalies and local importance scores vary when a significant feature such as xmeas_11 is paired with other pertinent attributes (i.e. xmeas_22 and xmeas_7) or with less salient ones (i.e. xmeas_12). In the first case, displayed in 2(b) and 2(a), anomalies (represented as red stars) exhibit a dispersed arrangement, resembling a bisector line, thereby affirming the considerable significance of both dimensions.
Moreover, the heatmap outlining the local importance scores partitions the feature space into two symmetric halves, indicating a shared importance between the respective pairs of features.
Conversely, in Figure 2(c) xmeas_11 is examined together with xmeas_12, one of the least important features in TEP, according to the GFI scores ranking. A notable distinction can be noticed in the distribution of inliers and outliers: anomalies align along the axis corresponding to feature xmeas_11, while normal points (illustrated as blue dots) are distributed along feature xmeas_12, underscoring the efficacy of this attribute in detecting normal points rather than anomalies. As a result, this visualization proves the reason why these two features are correctly placed at the two opposite ends of the GFI importance ranking by ExIFFI.
Finally, the predominance of blue coloring in the background heatmap signifies higher local importance scores for feature xmeas_12, indicative of anomalies being more readily discernible by exploiting vertical hyperplanes aligned with the direction of xmeas_11.
((a))
((b))
((c))
Figure 2: Local Scoremaps TEP dataset
4.4.3 Feature Selection Proxy Task
Utilizing labeled samples, different XAI models are compared on the TEP dataset via the Feature Selection proxy task, described in [10]. Two model-specific approaches (i.e. DIFFI [21] and ExIFFI) are considered alongside two model-agnostic algorithms: AcME-AD [23] and KernelSHAP [22].
All the models were employed to explain the $\text{EIF}^{+}$ AD model, except for DIFFI, tailored to the IF model. As described in 4.2.3 the proxy task requires producing a GFI ranking. However, AcME-AD and KernelSHAP algorithms can only produce LFI scores, so those are used for their rankings  [^4] For KernelSHAP 2% of the dataset is used as background and SHAP values are computed on the 100 most anomalous points.
Feature Selection plots are grouped in Figure 3. Comparing the $AUC_{FS}$ scores, the most effective explanations are produced by ExIFFI and AcME-AD. In these plots, the direct approach has a clearer decreasing trend compared to DIFFI and KernelSHAP, resulting in higher $AUC_{FS}$ values.
Figure 3: TEP feature selection for inverse (red), direct (blue), random (green) approaches.
Figure 4: CoffeeData feature selection for inverse (red), direct (blue), random (green) approaches.
4.5 Case Study II: PIADE Dataset
For the second case study, the PIADE dataset is considered. Differently from TEP, comprising synthetically generated data, PIADE incorporates alarm logs data coming from operating packaging machines lacking annotated samples. Consequently, the application of the Feature Selection proxy task is precluded, confining the experimental results to the assessment of Global and Local interpretability, as addressed in Sections 4.5.1 and 4.5.2.
4.5.1 Global Feature Importance
The lack of ground truth labels in PIADE requires tuning the dataset contamination (i.e. percentage of anomalous points) hyperparameter before computing the GFI scores. This hyperparameter is in fact needed to divide the data into inliers and outliers to compute the $\hat{I}_{O}$ and $\hat{I}_{I}$ terms, as detailed in 3. After careful tuning, the results presented in Figure 4.5.1 are obtained with a contamination factor of 15%.
Feature %scheduled_downtime stands on the top of the ranking. This variable quantifies the duration of downtime resulting from scheduled maintenance. It is reasonable to assume that if several dangerous alarms are triggered, resulting in serious damage, maintenance operations will be scheduled confirming the plausibility of this prediction. High importance scores can also be observed for A_010 and A_017 which are known machine failures according to domain experts [24].
4.5.2 Local Interpretability
Figure 5 gathers together three distinct Local Scoremaps outlining the relationship between variable %scheduled_downtime and various alarm-related features.
Taking inspiration from Section 4.4.2 multiple feature pairs are analyzed employing the Local Scoremaps to observe how variable %scheduled_downtime relates to different alarm related features.
Figure 5(a) juxtaposes the two most important features, as delineated by the Score plot in Figure 4.5.1: %scheduled_downtime and A_010. The scoremap reveals a clear L-shaped distribution formed by anomalous points (i.e. red stars), perfectly aligned with features %scheduled_downtime and A_010, confirming their pivotal role in the anomaly detection task. The distinctive shape of the anomalies distribution can be interpreted as follows: during scheduled downtime phases alarms are not triggered, resulting in anomalies scattered along the horizontal axis.
Conversely, during the production stage, alarms such as A_010 and A_017 are raised, as depicted by vertically aligned anomalies in both Figure 5(a) and 5(b). Among the aforementioned alarms, the dispersion of outliers is more pronounced in Figure 5(a), underscoring the decisive role of A_010 in distinguishing between normal and abnormal samples.
Finally, in Figure 5(c), %scheduled_downtime is related to the least important feature, as per the GFI score ranking produced by ExIFFI (i.e. %idle). The idle condition is the opposite operational state with respect to scheduled_downtime thus most inliers are aligned with the vertical dimension of the scoremap while the majority of the outliers are scattered along the axis representing %scheduled_downtime, as expected.
((a))
((b))
((c))
Figure 5: Local Scoremaps PIADE dataset
4.6 Case Study III: CoffeeData Dataset
This section reports the results of the experiments conducted on the CoffeeData dataset which distinguishes itself from the others for the time dependent nature of its data leading to the production of different plots to showcase the models’ performances.
4.6.1 Global Feature Importance
The assessment of global interpretability produced by ExIFFI on CoffeeData is realized through a different visualization whose aim is to highlight the temporal nature of the time-series data at our disposal.
The top half of Figure 6 reports the average time series divided into normal and anomalous data. In the bottom, the distribution of the GFI scores across multiple runs of each feature is represented with a box-plot, with the median highlighted by an orange line, which shows the time evolution of importance scores.
Figure 6: GFI Score Plot CoffeeData dataset.
Analyzing the GFI plot, a significant increase in the importance score can be observed between sample 8 and sample 11, which correspond to the highest difference between normal and anomalous time series, as observable in 6. This result confirms the ability of the ExIFFI interpretation to identify a good moment to signal to the user about an anomalous usage of the equipment.
4.6.2 Feature Selection Proxy Task
Figure 4 depicts the Feature Selection plots for the four interpretations under analysis. The plot is produced with the same settings employed in 4.4.3 and confirms the superiority of ExIFFI in ranking features in terms of importance with respect to other state-of-the-art approaches. Notably, AcME-AD and DIFFI produce a negative $AUC_{FS}$ score meaning that useless features were wrongly ranked in top positions.
4.7 Time Comparison Experiment
Time efficiency is a key requirement in the deployment of ML models in IIoT. Alarms described in PIADE may be triggered at a frequency of four alarms per minute [29]. Consequently, AD models should swiftly detect anomalies to avoid catastrophic events. Accordingly, in this section the time efficiency of ExIFFI is compared to the one of DIFFI, AcME-AD and KernelSHAP, introduced in 2.
The time comparison test assesses the time taken by each one of the models under examination to generate LFI explanation for a single anomalous point  [^5] The experiments were performed using an Intel i5 processor with 4 cores, 64 bit, 2.8 GHz, RAM 16 GB. Due to limited computational resources, KernelSHAP could only use sub-sampled versions of TEP and PIADE (2% and 25% respectively) as the background data used to fit the explainer.
Table 4 outlines the time performances of the four models considered on both datasets. Model-specific approaches (i.e. DIFFI and ExIFFI) exhibit efficient computational performance, while model-agnostic models (i.e. AcME-AD and KernelSHAP) demonstrate significantly lower efficiency.
In particular the high computational burden of KernelSHAP, makes its use impractical for industrial environments, considering it is using a sub-sampled version of the original dataset, which leads to inaccurate explanations.
Comparing the computational performances of the three datasets, PIADE exhibits higher time values. PIADE has an elevated feature count which highly affects the KernelSHAP and AcME-AD complexities. CoffeeData is the dataset with the lowest number of attributes in fact it is associated to the faster explanation times.
Conversely, model-specific interpretability models, DIFFI and ExIFFI, still present exceptional computation efficiency, independently on the number of features. The implementation of these models, laveraging the high-performing IF and EIF/ $\text{EIF}^{+}$, is the other reason under the speed up provided by these models. Their only drawback is the limited flexibility due to the fact that they are tailored to explain just a specific model (i.e. EIF/ $\text{EIF}^{+}$ and IF).
It is noteworthy that in all datasets ExIFFI is one order of magnitude faster than DIFFI thanks to the usage of the C programming language to optimize specific code segments using parallel computing.
Table 4: LFI explanation times (in seconds) for a single sample by ExIFFI, AcME-AD, DIFFI, and KernelSHAP on TEP, PIADE, and CoffeeData datasets.
Method TEP PIADE CoffeeData
ExIFFI 0.016 0.017 0.014
DIFFI 0.082 0.070 0.107
AcME-AD 3.60 16.5 0.523
KernelSHAP 112.42 138.94 4.99
5 Conclusions
We demonstrate the successful application of ExIFFI for industrial anomaly detection, showing how it fosters decision-making processes in real-world scenarios, providing insightful explanations of model predictions through informative visualizations, enabling the integration between efficient algorithmic performances and domain knowledge of expert operators.
ExIFFI was tested on three datasets coming from different industrial settings, equipped with high data dimensionality and unlabeled data points. The explanations provided by ExIFFI aligned with the ground truth on the anomalies’ root cause features. The effectiveness of the interpretations was proved against state-of-the-art interpretability methods through the Feature Selection proxy task, as shown in Sections 4.4.3,4.6.2.
We highlight the time efficiency of our method in detecting anomalies and providing explanations to foster Root Cause Analysis. The computational aspect is crucial in the industrial scenario where fast responses are required. ExIFFI outperforms other interpretability methods present in the literature in this aspect as outlined in 4.7. This result is achieved thanks to the nature of ExIFFI, which is a model-specific interpretability algorithm, paired to EIF, one of the most time-efficient models in the field of AD.
Future research includes the deployment of ExIFFI in an IIoT system for real-time industrial equipment AD and feature importance. The computational efficiency and low memory needs of ExIFFI can be applied in tiny ML [31], deploying the model in devices with limited computational and memory resources such as Micro Controller Units (MCUs) commonly used in IIoT.