### v3

# Davide Frizzo

Abstract
Anomaly Detection (AD) is crucial in industrial settings to streamline operations by detecting underlying issues. Conventional methods merely label observations as normal or anomalous, lacking crucial insights. In Industry 5.0, interpretable outcomes become desirable to enable users to understand the rational under model decisions. This paper presents the first industrial application of ExIFFI, a recent approach for fast, efficient explanations for the Extended Isolation Forest (EIF) AD method. ExIFFI is tested on three industrial datasets, demonstrating superior explanation effectiveness, computational efficiency and improved raw anomaly detection performances. ExIFFI reaches over then 90% of average precision on all the benchmarks considered in the study and overperforms state-of-the-art Explainable Artificial Intelligence (XAI) approaches in terms of the feature selection proxy task metric which was specifically introduced to quantitatively evaluate model explanations.
Keywords Anomaly Detection, Explainable Artificial Intelligence, Industrial Internet of Things
AI
Artificial Intelligence
AD
Anomaly Detection
IF
Isolation Forest
EIF
Extended Isolation Forest
DIFFI
Depth Based Isolation Forest Feature Importance
AcME-AD
Accelerated Model Explanations for Anomaly Detection
ML
Machine Learning
IoT
Internet of Things
IIoT
Industrial Internet of Things
XAI
Explainable Artificial Intelligence
SHAP
SHapley Additive exPlanations
GFI
Global Feature Importance
LFI
Local Feature Importance
PIADE
Packaging Industry Anomaly DEtection
TE
Tennessee Eastman
TEP
Tennessee Eastman Process
LODA
Lightweight on-line detector of anomalies
1 Introduction
In recent decades, the rapid expansion of Machine Learning (ML) applications has significantly propelled advancements in Internet of Things (IoT) technologies, leading to the widespread deployment of interconnected sensor networks [16]. These developments are particularly notable in the context of Industrial Internet of Things (IIoT), which focuses on real-time monitoring of industrial assets to optimize maintenance and enhance production efficiency [23]. As these IoT systems generate vast datasets, ML algorithms play a crucial role in analyzing this data to extract predictive insights, underscoring a symbiotic relationship between ML and IIoT that mutually enhance their capabilities [3].
A key application within this context is unsupervised AD [12], especially prominent in settings where labeling data is impractical. Among AD methods, models based on Isolation Forest (IF[19], in particular EIF [17] stands out due to its speed, low memory requirements, and high performance, making it valuable in industrial settings for swiftly pinpointing anomalies with exceptional performances [4]. However, the challenge extends beyond merely detecting anomalies; understanding their causes is essential for effective resolution [9]. Here, XAI becomes crucial by making the outputs of ML models like EIF transparent and actionable, thus supporting informed decision-making and enhancing system reliability within IIoT environments [1].
This paper specifically evaluates the Extended Isolation Forest Feature Importance (ExIFFI) algorithm, developed by Arcudi et al. [2], which provides a time-efficient and tailored interpretative approach to the EIF model. Moreover, [2] additionally introduces a modification of the EIF algorithm, named Extended Isolation Forest Plus ( $\text{EIF}^{+}$) which improves the generalization performances of EIF and can also be interpreted by ExIFFI. The algorithm’s effectiveness is showcased through its application on real world IIoT datasets, demonstrating how it enhances the decision-making process in industrial settings.
This paper is the extension of [15]. The improvements over the original version focus on providing additional, refined and more robust experimental results to showcase the applicability of EIF, $\text{EIF}^{+}$ and ExIFFI in industrial environments. In particular they consists of:
  • An evaluation of the explained models in terms of AD performances, detailed in 4.3.
  • ExIFFI performance assesment for the Packaging Industry Anomaly DEtection (PIADE) dataset is extended to consider data from all the five machines included in the study as presented in 4.5.1.
  • The effect of hyperparameter such as number of trees and contamination factor on ExIFFI performances is assesed through ablation studies showcased in 4.8.
  • Additional images to showcase the results (i.e. Figures 2(a), 2(b), 4).
  • The addition of a new benchmark dataset, CoffeeData, introduced in 4.1.1.
The paper is organized as follows: in Section 2 ExIFFI is contextualized within Industry 5.0, Section 3 explains the operational mechanics of the algorithm, Section 4 presents the benchmark datasets and the application of ExIFFI on them and ablation studies are reported in 4.8. Finally, Section 5 summarizes the key findings and discusses potential directions for future research.
2 Related Work
The progression from Industry 4.0 [18] to Industry 5.0 [26] marks a shift from a focus on enhancing operational performance through automation and data exchange, primarily using cyber-physical systems and IoT, to emphasizing human-centric outcomes [31]. Industry 5.0 integrates human creativity with the technical prowess of Artificial Intelligence (AI) and robotics, highlighting the need for transparency and interpretability in machine learning models to build trust and aid decision-making.
This need is addressed by implementing XAI in various industrial applications, including fault detection in machinery [5], process monitoring across sectors from semiconductors to home appliances [7, 14, 8], and predictive maintenance [27]. The ExIFFI interpretation algorithm of the EIF model used for AD exemplifies this trend. It offers a fast, precise interpretation method crucial for IIoT environments, where performance and user-centric understanding are imperative for efficiency and accessibility in modern manufacturing setups.
In this paper ExIFFI is evaluated against various interpretation techniques, including both ad-hoc  [^1] Ad-hoc interpretation algorithms are built into models for inherent transparency. Post-hoc methods clarify complex model decisions after development but can be computationally intensive and approximate. [21] methods like Depth Based Isolation Forest Feature Importance (DIFFI[10], tailored for the Isolation Forest, and post-hoc methods such as KernelSHAP [20] and Accelerated Model Explanations for Anomaly Detection (AcME-AD[29, 30].
KernelSHAP is notable for its ability to interpret diverse models through a model-agnostic, kernel-based approximation of SHapley Additive exPlanations (SHAP) values. However, it can be computationally demanding, particularly with large datasets. In contrast, AcME-AD streamlines the explanation process, making it ideal for scenarios that require swift interpretability compared to KernelSHAP.
Although ExIFFI and DIFFI are less flexible compared to post-hoc methods, they significantly boost computational speed, delivering reliable interpretations up to 100 times faster than post-hoc models as shown in Table 6. Additionally, ExIFFI leverages the advanced structure of the EIF, that outperforms the traditional IF [4] by avoiding its structural biases [17]. This enables ExIFFI to assess feature significance both efficiently and accurately.
3 Proposed Approach
The ExIFFI, similarly to the DIFFI with the IF, leverages the structure of the forest generated by the EIF to assess the significance of each feature in determining whether a sample is anomalous. As a consequence ExIFFI can be thought as a wrapper to EIF and $\text{EIF}^{+}$ that adds interpretability over the model’s outputs without affecting the AD performances by any means. The importance computation is in fact naturally embedded in the model fitting procedure.
Note that, as in IF, the EIF method hypothesizes that anomalous points are located in low density regions of the input space and thus are easy to isolate from inliers. For these reasons the algorithm focuses on identifying isolated anomalies and it’s not able to identify subtle variations within dense clusters like [28].
In the structure of the EIF, each tree $t$ in the forest $\mathcal{T}$ consists of nodes that partition the space using a hyperplane $\mathcal{H}_{k}^{t}$. Each hyperplane is defined by a normal vector $\mathbf{v}_{k}^{t}$ and an intercept point $\mathbf{p}^{t}_{k}$. The hyperplane splits the data sample $X_{k}^{t}$ into two distinct subsets $L^{t}_{k}$ and $R^{t}_{k}$, with $L^{t}_{k}\cup R^{t}_{k}=X_{k}^{t}$ and $L^{t}_{k}\cap R^{t}_{k}=\varnothing$. The ExIFFI algorithm assesses the feature importance by calculating the imbalance generated by each node for a given sample $x$ as follows:
$\mathbf{\lambda}^{t}_{k}(\mathbf{x})=\begin{cases}\left(\frac{|X^{t}_{k}|}{|L^{t}_{k}|}\right)\text{abs}(\mathbf{v}^{t}_{k}),&\text{if }\mathbf{v}^{t}_{k}\cdot\mathbf{x}>\mathbf{v}^{t}_{k}\cdot\mathbf{p}^{t}_{k}\\ \left(\frac{|X^{t}_{k}|}{|R^{t}_{k}|}\right)\text{abs}(\mathbf{v}^{t}_{k}),&\text{otherwise}\end{cases}$ (1)
Differently from IF, the non axis-aligned hyperplanes of EIF help to better detect anomalies spread along multiple dimensions. Consequently explainability is enhanced even in highly dimensional datasets with importance scores naturally distributed among multiple important features.
The feature importance for a sample $x$ within a specific tree $t$ is determined by the summation of importance vectors from all nodes $k$ that $x$ traverses on its path to the leaf:
$\mathbf{I}_{t}(x)=\sum_{k\in\mathcal{P}^{t}_{x}}\mathbf{\lambda}_{k}^{t}(x)$ (2)
The overall feature importance for $x$ across the entire forest is then computed by aggregating these importance vectors from all the trees that $x$ encounters:
$\mathbf{I}(x)=\sum_{t\in\mathcal{T}}\mathbf{I}_{t}(x)$ (3)
This importance vector $\mathbf{I}(x)$ assesses the influence of each feature on the isolation of $x$, identifying the ones predominantly contributing to the classification $x$ as an anomaly. To correct for potential biases caused by features being sampled more frequently, $\mathbf{I}(x)$ is normalized by the sum of vectors orthogonal to the hyperplanes of the nodes that $x$ traverses in each tree, denoted as:
$\mathbf{V}(x)=\sum_{t\in\mathcal{T}}\sum_{k\in\mathcal{P}^{t}}\mathbf{v}_{k}^{t}$ (4)
Grouping together the importance vectors assigned to all the samples in the input dataset, ExIFFI is equipped with two possible outputs: Global Feature Importance (GFI) and Local Feature Importance (LFI). GFI returns a single vector which associates a score to each feature, quantifying its overall importance in discriminating between inliers and outliers. It can be obtained as follows:
$\mathbf{GFI}=\frac{\hat{\mathbf{I}}_{O}}{\hat{\mathbf{I}}_{I}}$ (5)
where $\hat{\mathbf{I}}_{O}$ and $\hat{\mathbf{I}}_{I}$ are the importance vectors computed over the set of outliers and inliers respectively. On the other hand, the Local Feature Importance ($\mathbf{LFI}$) for $x$ is calculated as:
$\mathbf{LFI}(x)=\frac{\mathbf{I}(x)}{\mathbf{V}(x)}$ (6)
This metric offers a refined and normalized measure of which features significantly impact the classification of a single sample $x$ as anomalous. In this way the interpretability of AD models is enhanced and facilitates targeted interventions based on the identified critial features.
4 Experimental Results
This section presents the results of applying ExIFFI to two publicly accessible and a private dataset derived from industrial processes, which serve as benchmarks for evaluating ExIFFI’s effectiveness within real-world contexts. The datasets include Tennessee Eastman Process (TEP), which offers synthetic data with established ground truth for anomaly-inducing features [22], PIADE, which encapsulates typical real-world challenges such as unlabeled and high-dimensional data [25]. Finally, CoffeeData is a confidential dataset containing sensor measurements recorded during the brewing phase of capsules on coffee machines.
The section is organized as follows: firstly, datasets are presented in 4.1, the experimental setup is provided in 4.2, AD model performances are reported in 4.3, global interpretability is assessed in 4.4.1, 4.6.1 respectively, and four different interpretability models are compared in 4.4.2, 4.6.2 by means of the Feature Selection Proxy Task.
The outcomes here discussed are achieved using Python as the base language to implement the method and C to optimize functions embedded within the performance-critical segments of the Python code. [^2] The source code of this project is available in a public repository, with reproducible results: https://github.com/francesco-borsatti-unipd/ExIFFI_Industrial_Test.
4.1 Industrial IoT datasets
In the following, we provide a description of the structure of the benchmark datasets used in this study (i.e. TEP, PIADE and CoffeeData) detailing their key characteristics and differences.
The datasets exhibit fundamental disparities representing different industrial scenarios. Specifically PIADE comprises machine alarms and statuses, while TEP and CoffeeData are characterized by time-series data. These differences underscore the wide range of signal types on which the presented approach can be applied. Another key difference between the three relies on how data are generated: PIADE and CoffeeData are real datasets collected from existing machines, while TEP was obtained through simulations (albeit precise).
Table 1: The main characteristics of the industrial datasets used for this study are reported: dataset size, sampling rate, source and presence of labels
Dataset Samples Features
Sampling
Period $[s]$
Anomaly
Labels
Source
PIADE 23376 165 3600 No [25]
TEP 35600 52 180 Yes [22]
CoffeeData D1 372 15 0.5 Yes confidential
CoffeeData D2 255 15 0.5 Yes confidential
A summary of the main datasets characteristics is reported in Table 1.
For a more detailed description of the TEP and PIADE datasets please refer to [15].
4.1.1 CoffeeData Dataset
The CoffeeData dataset was firstly introduced in [24]. It is a confidential dataset and consists of multiple time-series representing the brewing process of capsules in coffee machines. The aim of AD in this scenario is to swiftly identify anomalous usages of coffee capsules in order to improve the brewing quality. Anomalous behaviors include using the same capsule twice or more times (i.e. reused capsule) or using the coffee machine without any capsule inside it (i.e. no capsule).
The data were collected through a Yokogawa MX100 device using a sampling frequency of 2 Hz. Eight variables are registered including: heating element voltage and current, water flow rate, temperatures for the heating element bulk, output water, ambient air, water tank, and water pump.
Following the procedure introduced in [24], only the water flow rate feature was considered, as it was the most informative one.
Each time series is trimmed to the first 15 samples of the pre-infusion phase, as this segment provides enough information to address the problem effectively while also allowing for early detection of anomalies, which is critical for practical use.
In order to take into account the temporal dimension embedded in the data, instead of employing tedious and time-consuming handcrafted features as in [24], raw data are used and time stamps are considered as the features which will be ranked by ExIFFI when computing importance scores. Exploiting this strategy is possible to identify the time instants in which anomalies occur. A possible application of this result may be to trigger an alarm at the anomalous time sample to notify the user of an improper usage of the equipment.
Table 2 outlines the dataset composition: standard capsules are part of the inliers group (comprising both original and compatible brands), while there are two kinds of anomalous samples: reused capsules and no capsules.
Table 2: CoffeeData dataset composition.
Brewing Capsule # Samples Class
1st standard 39 Normal
2nd standard 229
any reused capsule 78 Anomaly
no capsule 26
Normal capsules are divided into the first and second brewings categories:
  • First Brewing: the coffee machine is used for the first time after a long period, so it has to reach the working temperature
  • Second Brewing: multiple coffees are produced in a short time interval so the machine is already at the working temperature
Clearly, trials of type second brewing are easier to carry out (i.e. there is no need to wait for the machine to cool down) and thus the amount of acquired samples in this condition is higher.
4.2 Experimental Setup
This section provides a description of the techniques used to evaluate and compare the different interpretability methods considered in this study.
In order to provide both a qualitative and quantitative assesment of the ExIFFI performances three different kind of experiments are conducted:
  • Global Importance Assessment: This experiment produces a Score Plot ranking the features in decreasing order of GFI score. It represents a powerful tool for human operators to discover the root causes of anomalies and take preventive actions.
  • Local Scoremaps: Local scoremaps focus on local interpretability by selecting a pair of features and showing how inliers and outliers are distributed along them. An heatmap is superimposed to showcase the distribution of the LFI scores in the feature space.
  • Feature Selection Proxy Task: In this study feature selection is used as a proxy task to provide a quantitative measure of the effectiveness of a XAI approach by plotting the trend of the average precision metric as features are progressively removed in increasing or decreasing order of GFI scores.
In order to take into account the stochasticity of ExIFFI, in all the above mentioned experiments multiple runs with different seeds were conducted in order to provide robust results and showcase their statistical significance.
For the sake of space results on the Local Scoremaps experiments are not included in this extension. The interested reader can find a detailed analysis of local interpretability in [15].
4.3 AD Performance Comparison
In this section IF, EIF and $\text{EIF}^{+}$ AD models are compared with typical AD metrics for the labeled datasets (i.e. TEP and CoffeeData). Tables 3, 4 report the test set metrics. This evaluation is a crucial step for the application of XAI methods like ExIFFI since their effectiveness highly relies on the accuracy of the explained model.
Table 3: TEP dataset AD metrics for the considered models. Best performance for each metric in bold.
Model Average Precision Precision ROC AUC Fit Time [s] Predict Time [s]
IF 0.90 0.85 0.92 1.90 1.44e-4
EIF 0.93 0.90 0.95 1.25 8.00e-5
$\text{EIF}^{+}$ 0.93 0.91 0.95 1.71 5.14e-5
Table 3 shows how EIF and $\text{EIF}^{+}$ perform better than IF thanks to their higher-complexity data partitioning scheme, leveraging multidimensional splits as opposed to IF.
CoffeeData comprises two evaluation datasets: D1, which contains both first and second brewings and both kinds of anomalies (i.e. no capsule and reused capsule); while D2 is more limited, but equivalent to the one used in previous work [24] (i.e. only second brewing data and no capsule anomalies).
Table 4: CoffeeData dataset AD metrics for the considered models. Best performance for each metric in bold.
Dataset Model Average Precision Precision ROC AUC Fit Time [s] Predict Time [s]
D1 IF 0.74 0.65 0.63 1.18 5.48e-5
EIF 0.77 0.66 0.63 0.81 3.65e-5
$\text{EIF}^{+}$ 0.79 0.65 0.64 0.99 4.87e-5
D2 IF 0.91 0.96 0.97 0.96 3.04e-5
EIF 0.97 0.89 0.98 0.74 5.48e-5
$\text{EIF}^{+}$ 0.99 0.96 0.99 0.96 4.26e-5
AD model performances are showcased in Table 4. The increased difficulty associated with D1 translates in slightly worse AD metrics, but still the best performing model is $\text{EIF}^{+}$ followed by EIF and IF. For what concerns D2 metrics are on par with the ones showcased in [24] where the model was exploiting hand crafted features while raw time samples are used in this setting.
4.4 Case Study I: TEP Dataset
4.4.1 Global Feature Importance
Figure 1 exhibits the Score Plot for the TEP dataset. Features xmeas_22 and xmeas_11 emerge as more relevant than others, aligning with the ground truth [13] since xmeas_11 is the root cause for fault IDV12. Furthermore, the Sign Directed Graph in [13] proves also the saliency of xmeas_22 (i.e. the Separator cooling water outler temperature), a direct consequence of xmeas_11. The causal relationship between these two attributes leads the model to position them on top of the GFI ranking.
Figure 1: This plot represents the top 8 features of the GFI ranking returned by ExIFFI on TEP.
4.4.2 Feature Selection Proxy Task
Utilizing labelled samples, different XAI models are quantitatively compared on the TEP dataset via the Feature Selection proxy task, described in [2]. Two model-specific approaches (i.e. DIFFI [10] and ExIFFI) are considered alongside two model-agnostic algorithms: AcME-AD [29] and KernelSHAP [20].
All the models were employed to explain the $\text{EIF}^{+}$ AD model, except for DIFFI, tailored to the IF model. As described in 4.2 the proxy task requires producing a GFI ranking. However, AcME-AD and KernelSHAP algorithms can only produce LFI scores, so those are used for their rankings  [^3] For KernelSHAP 2% of the dataset is used as background and SHAP values are computed on the 100 most anomalous points.
Feature Selection plots are grouped in Figure 2(a). Comparing the $AUC_{FS}$ scores, the most effective explanations are produced by ExIFFI and AcME-AD. In these plots, the direct approach has a clearer decreasing trend compared to DIFFI and KernelSHAP, resulting in higher $AUC_{FS}$ values. KernelSHAP struggles to rank features effectively, due to the limitations imposed by the sub-sampled background data to make it computationally feasible, leading to a negative $AUC_{FS}$ score. Moreover the standard deviation of the average precision values is in the order of $10^{-3}$ confirming the statistical significance of the results.
((a)) TEP dataset
((b)) CoffeeData dataset
Figure 2: Feature selection results for inverse (red), direct (blue), and random (green) approaches using DIFFI, AcME-AD, KernelSHAP, and ExIFFI XAI: (a) TEP, (b) CoffeeData.
4.5 Case Study II: PIADE Dataset
For the second case study, the PIADE dataset is considered. Differently from TEP, comprising synthetically generated data, PIADE incorporates alarm logs data coming from operating packaging machines lacking annotated samples. Consequently, the application of the Feature Selection proxy task is precluded, confining the experimental results to the assessment of Global interpretability, as addressed in Sections 4.5.1.
4.5.1 Global Feature Importance
The lack of ground truth labels in PIADE requires tuning the dataset contamination (i.e. percentage of anomalous points) hyperparameter before computing the GFI scores. This hyperparameter is in fact needed to divide the data into inliers and outliers to compute the $\hat{I}_{O}$ and $\hat{I}_{I}$ terms, as detailed in 3, and can have a crucial impact on AD and XAI performances as it will be shown in 4.8.2 and 4.8.3. After careful tuning, the results presented in Figure 3 are obtained with a contamination factor of 15%.
Feature %scheduled_downtime stands on the top of the ranking. This variable quantifies the duration of downtime resulting from scheduled maintenance. It is reasonable to assume that if several dangerous alarms are triggered, resulting in serious damage, maintenance operations will be scheduled confirming the plausibility of this prediction. High importance scores can also be observed for A_010 and A_017 which are known machine failures according to domain experts [30].
Figure 3: This plot represents the top 8 features of the GFI ranking returned by ExIFFI on PIADE.
Table 5: Top three features according to the GFI rankings produced in the different PIADE machines datasets
Dataset $1^{st}$ $2^{nd}$ $3^{rd}$
piade_s1 %idle A_066 %scheduled_downtime
piade_s2 %scheduled_downtime A_010 A_005
piade_s3 %idle %downtime A_008
piade_s4 idle/idle %idle A_010
piade_s5 %idle idle/idle A_005
Table 5 reports the top 3 features obtained from the rankings induced by GFI scores computed on the data from the 5 packaging machines included in PIADE. Taking into account that these machines are of the same type but are inserted in different working and environmental conditions the result show an high consistency in detecting features regarding idle and downtime intervals as important descriptors. On the other hand the differences in the operational environment of the machines introduces some variability in the specific alarm features included in the top 3 rankings.
4.6 Case Study III: CoffeeData Dataset
This section reports the results of the experiments conducted on the CoffeeData dataset which distinguishes itself from the others for the time dependent nature of its data leading to the production of different plots to showcase the models’ performances.
4.6.1 Global Feature Importance
The assessment of global interpretability produced by ExIFFI on CoffeeData is realized through a different visualization whose aim is to highlight the temporal nature of the time-series data at our disposal.
The top half of Figure 4 reports the average time series divided into normal and anomalous data. In the bottom, the distribution of the GFI scores across multiple runs of each feature is represented with a box-plot, with the median highlighted by an orange line, which shows the time evolution of importance scores.
Figure 4: The trend of the GFI scores increases in the time interval where there is the most significant difference between the time series of normal and anomalous capsules
Analyzing the GFI plot, a significant increase in the importance score can be observed between sample 8 and sample 11, which correspond to the highest difference between normal and anomalous time series, as observable in 4. This result confirms the ability of the ExIFFI interpretation to identify a good moment to signal the user about an anomalous usage of the equipment.
4.6.2 Feature Selection Proxy Task
Figure 2(b) depicts the Feature Selection plots for the four interpretations under analysis. The plot is produced with the same settings employed in 4.4.2 and confirms the superiority of ExIFFI in ranking features in terms of importance with respect to other state-of-the-art approaches. Notably, AcME-AD and DIFFI produce a negative $AUC_{FS}$ score meaning that useless features were wrongly ranked in top positions. Similarly to TEP also in this case the low variation of the average precision values over different runs proves that the results are statistically significant.
4.7 Time Comparison Experiment
Time efficiency is a key requirement in the deployment of ML models in IIoT. Alarms described in PIADE may be triggered at a frequency of four alarms per minute [11]. Consequently, AD models should swiftly detect anomalies to avoid catastrophic events. Accordingly, in this section the time efficiency of ExIFFI is compared to the one of DIFFI, AcME-AD and KernelSHAP, introduced in 2.
The experiment assesses the time taken by each one of the models under examination to generate LFI explanation for a single anomalous point [^4] The experiments were performed using an Intel i5 processor with 4 cores, 64 bit, 2.8 GHz, RAM 16 GB. Due to limited computational resources, KernelSHAP could only use sub-sampled versions of TEP and PIADE (2% and 25% respectively) as the background data used to fit the explainer.
Table 6 outlines the time performances of the four models considered on both datasets. Model-specific approaches (i.e. DIFFI and ExIFFI) exhibit efficient computational performance, while model-agnostic models (i.e. AcME-AD and KernelSHAP) demonstrate significantly lower efficiency.
In particular the high computational burden of KernelSHAP, makes its use impractical for industrial environments, considering it is using a sub-sampled version of the original dataset, which leads to inaccurate explanations.
Comparing the computational performances of the three datasets, PIADE exhibits higher time values: its high feature count highly affects the KernelSHAP and AcME-AD complexities. CoffeeData is the dataset with the lowest number of attributes and it is in fact associated to the faster explanation times.
Conversely, model-specific interpretability models, DIFFI and ExIFFI, still present exceptional computation efficiency, independently on the number of features. The implementation of these models, laveraging the high-performing IF and EIF/ $\text{EIF}^{+}$, is the other reason under the speed up in importance computation. Moreover the fact that each isolation tree focuses on a bootstrap sample, usually of limited size, of the data makes them efficient also on large scale datasets as the ones included in this study. Their only drawback is the limited flexibility due to the fact that they are tailored to explain just a specific model (i.e. EIF/ $\text{EIF}^{+}$ and IF).
It is noteworthy that in all datasets ExIFFI is one order of magnitude faster than DIFFI thanks to the usage of the C programming language to optimize specific code segments using parallel computing.
Table 6: LFI explanation times (in seconds) for a single sample by ExIFFI, AcME-AD, DIFFI, and KernelSHAP on TEP, PIADE, and CoffeeData datasets.
Method TEP PIADE CoffeeData
ExIFFI 0.016 0.017 0.014
DIFFI 0.082 0.070 0.107
AcME-AD 3.60 16.5 0.523
KernelSHAP 112.42 138.94 4.99
4.8 Ablation Studies
In this section ablation studies are performed on some key EIF and ExIFFI hyperparameters to assess their influence on model performances. The considered hyperparameters are the number of isolation trees used to fit the ensemble and the contamination factor (i.e. the percentage of anomalies in the data). The latter has a crucial importance in industrial settings where labelled data are not available and thus the contamination factor has to be set without any objective grounding.
In these experiments we observed the trend of some AD and interpretability metrics for different hyperparameters values. For this reason these were conducted on datasets containing ground truth knowledge on the anomalous data like TEP and CoffeeData.
For the sake of space only the results on the TEP dataset are reported but the ones obtained on CoffeeData are not much different.
4.8.1 Number of Trees
This experiment consists in tracking the variation of the average precision AD metric as the number of trees used to fit the underlying AD model increases. The higher the number of trees the higher the variability of data observed by the model which becomes more specialized in the task of detecting anomalous patterns. As a result average precision is expected to increase with the number of estimators in the ensemble. On the other hand additional computations are required so augmented fitting and prediction times are also expected.
Figure 5: The average precision metric increases as the number of trees used to fit the AD model increases
In Figure 5 it’s possible to observe an improvement in the average precision metric passing from 10 to 30 trees, while from 50 trees onwards the model performances saturate at a constant average precision of about 93%. This result suggests the fact that 50 trees are sufficient to perform optimally in the AD task for the TEP dataset. For the sake of space we do not report the graphical results related to the relation between computational time and number of trees but these both show a linear increasing trend with the number of estimators.
4.8.2 Contamination Prediction
In AD models the contamination factor is used to set a threshold on the anomaly scores in order to convert them into binary label to classify a point as anomalous or not. Consequently is crucial to correctly set this quantity, especially in industrial settings.
In order to study this effect we designed an ablation study where different contamination levels are used to perform prediction with the $\text{EIF}^{+}$ model and the ROC AUC metric is considered for the evaluation of the AD performances. Contamination values both higher and lower then the true dataset contamination (i.e. about 4% for TEP) are considered. The aim is to evaluate the effects of false negative and false positive errors that are committed when the contamination is underestimated or overestimated, respectively.
Figure 6: Impact of contamination level on the ROC AUC score (top figure) and on the $AUC_{FS}$ metric. In both cases the best performances are obtained near the true dataset contamination value, indicated by the red vertical line. The x-axis is visualized in logarithmic scale.
The top plot of Figure 6 shows the impact of contamination level on ROC AUC performance, exhibiting a bell-shaped curve whose maximum occurs at approximately the true dataset contamination value. Notably the curve is almost symmetric, meaning that a significant overestimation produces a similar effect on the ROC AUC score as an underestimation.
4.8.3 Contamination Feature Selection
In the final ablation study we focus on interpretability by considering the $AUC_{FS}$ metric introduced in Section 4.2 to quantitatively evaluate the feature rankings produces by a XAI model. Also in this case contamination is considered as the varying hyperparameter since it has an effect also in the GFI computation as explored in Section 3. The same contamination values employed in Section 4.8.2 are considered.
The result is contained in the bottom part of Figure 6. Similarly to the upper part an U shaped curve can be observed. However the $AUC_{FS}$ values for small contamination factors are much smaller than the ones obtained for overestimated contaminations. This is probably due to the fact that the model has only a few anomalies to exploit for the computation fo the GFI scores in case the contamination is highly underestimated. This result is in accordance to the fact that it’s preferrable to overestimate the amount of anomalous samples in the dataset since false positive errors have less potentially dangerous consequences than false negative ones.
5 Conclusions
We demonstrate the successful application of ExIFFI for industrial anomaly detection, showing how it fosters decision-making processes in real-world scenarios. This is achieved by providing insightful explanations of model predictions through informative visualizations, enabling the integration between efficient algorithmic performances and domain knowledge of expert operators.
ExIFFI was tested on three datasets coming from different industrial settings, equipped with high data dimensionality and unlabeled data points. The explanations provided by ExIFFI aligned with the ground truth on the anomalies’ root cause features. The effectiveness of the interpretations was proved against state-of-the-art interpretability methods through the Feature Selection proxy task, as shown in Sections 4.4.2,4.6.2.
We highlight the time efficiency of our method in detecting anomalies and providing explanations to foster Root Cause Analysis. The computational aspect is crucial in the industrial scenario where fast responses are required. ExIFFI outperforms other interpretability methods present in the literature in this aspect as outlined in 4.7. This result is achieved thanks to the nature of ExIFFI, which is a model-specific interpretability algorithm, paired to EIF, one of the most time-efficient models in the field of AD.
On the other hand the limitations of the presented approach are related to its limited flexibility being ExIFFI tailored to explain only the EIF AD model. Another potential challenge to consider is the inability of such AD model to detect clustered anomalies.
For what concerns the industrial deployment of ExIFFI, its computational efficiency and low memory requirements makes it eligible for the application in a tiny ML [6] framework part of an IIoT infrastructure. Thus ExIFFI fits perfectly as a real-time industrial AD and feature importance solution in the context of Industry 5.0 which directs its focus towards human-aligned industrial equipments.